OpenStudy (anonymous):

A glove and a ball together costs $85. If the glove glove costs $75 more than the ball, what is the cost of the ball?

OpenStudy (anonymous):

$85 total, the glove costs 75$, so 85-75=10...the ball costs $10, hope that helps! :)

OpenStudy (anonymous):

k, thats what i had originally thought, thanks for the clarification !:D

OpenStudy (anonymous):

wait that's wrog!

OpenStudy (anonymous):

wrong*

OpenStudy (anonymous):

itsn't $5

OpenStudy (anonymous):

the correct response?

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

sorry I messed that up, haha

OpenStudy (anonymous):

k, well if one costs $10 and the other costs $75, the the glove costs $85, so then $85 +10 would equal 95, not 85

OpenStudy (anonymous):

k, well i figured it out, thanks for the effort anyways. im not fanning you btw

OpenStudy (anonymous):

oh, sorry, I misread the problem. But in any case, you understnad what you're doing, which is great!

OpenStudy (anonymous):

I don't care, sorry for not reading it right

OpenStudy (anonymous):

(: thanks anyway!

OpenStudy (anonymous):

can you put it in an equation to demonstrate how the anser is 5? please?

OpenStudy (anonymous):

I'll try...g=cost of glove....b=cost of ball..... (1) g + b=85 (2) g=75 + b ....then you plug 2 into 1

OpenStudy (anonymous):

Let x be the cost of the ball and y be the cost of the glove. x+y=85 x+75=y Solve for x

OpenStudy (anonymous):

how would that equal 5?

OpenStudy (anonymous):

75 + b +b=85..... 2b=85-75.... b=10/2.... b=5

OpenStudy (anonymous):

ok, but why did we have 2b?

OpenStudy (anonymous):

b + b = 2b...it's like saying 1b+ 1b

OpenStudy (anonymous):

ya ya, but why didwe get 2b, i mean wy didwe ad b + b

OpenStudy (anonymous):

ok we established that g=75+b, so we plugged that into the g+b=85 equation...that gave us (75+b)+b=85 ... if you have the cost of one glove + the cost of one glove you have $5 + $5, which equals 10$... 2(b) will also give us 10 because 2*5=10.

OpenStudy (anonymous):

it makes mathematical sense to represent the information that way

OpenStudy (anonymous):

by the way, g=75+b is an equation that means "the glove costs $75 more than the ball"

OpenStudy (anonymous):

ok thank you sooo much, i'm sorry about condesending you earlier, :DDDD

OpenStudy (anonymous):

No problem, glad I could help this time haha